# Mortgage Burnout and Selection Effects in Heterogeneous Cox Hazard Models

### 12 Mar 2026 v1

@@ 2 Mortgage Pool Heterogeneity
Let $(\Omega,{\mathscr{F}},({\mathscr{F}}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space describing the time evolution of the pool of mortgages, and let $(X,{\mathcal{X}},F)$ be a measurable type space representing the unobserved borrower heterogeneity. For each $x\in X$, let $\lambda_{t}(x)$ be a nonnegative ${\mathscr{F}}_{t}$-adapted intensity process. Typically, $\lambda_{t}$ depends explicitly on a number of loan characteristics $y$, such as refinance incentive (CM(t))(C-M(t)), size, age, loan to value ratio, FICO score, geographic region, which represent the observed heterogeneity. Unless stated otherwise, we will suppress this dependence for the sake of notational convenience.
@@
Define the survival weighted cross sectional measure
@@ 4.1 Common Factor Hazard Dynamics
Since the pool hazard satisfies
@@ 4.2 Measure Change Interpretation
This representation shows that burnout can be interpreted as a measure change phenomenon: the observed pool dynamics correspond to expectations taken under a survival biased distribution of borrower types. In other words, the cross sectional distribution of borrower types observed in a seasoned mortgage pool is endogenously tilted toward lower hazard borrowers. Burnout therefore arises as a purely statistical selection effect rather than a behavioral change in borrower refinancing incentives.
@@ 5.3 Normal Frailty
For illustration we also consider the normal distribution,
@@ 6 Conclusion
The negative variance term provides a structural stochastic analog of burnout: heterogeneity induces a systematic downward drift in the observed hazard through selection effects. The resulting formula may be interpreted as a stochastic continuous-time analog of the Price equation, showing that burnout arises from a universal selection mechanism in heterogeneous hazard populations.

### 16 Mar 2026 v2

@@ 2 Mortgage Pool Heterogeneity
Let $(\Omega,{\mathscr{F}},({\mathscr{F}}_{t})_{t\geq 0},\mathsf{P})$ be a filtered probability space describing the time evolution of the pool of mortgages, and let $(X,{\mathscr{X}},F)$ be a measurable type space representing the unobserved borrower heterogeneity. For each $x\in X$, let $\lambda_{t}(x)$ be a nonnegative ${\mathscr{F}}_{t}$-adapted intensity process. Typically, $\lambda_{t}$ depends explicitly on a number of loan characteristics $y$, such as refinance incentive (the difference between the coupon on the loan and the current mortgage rate), size, age, loan to value ratio, FICO score, geographic region, which represent the observed heterogeneity. Unless stated otherwise, we will suppress this dependence for the sake of notational convenience.
@@
Define the survival weighted cross-sectional measure
@@ 4.1 Common Factor Hazard Dynamics
Idiosyncratic borrower shocks diversify in large pools and therefore do not contribute to the aggregate hazard dynamics. Since the pool hazard satisfies
@@
Similar selection effects arise in heterogeneous credit hazard models, where dispersion in firm level default intensities leads to declining average hazard among surviving firms.
@@ 4.2 Measure Change Interpretation
This representation shows that burnout can be interpreted as a measure change phenomenon: the observed pool dynamics correspond to expectations taken under a survival biased distribution of borrower types. In other words, the cross-sectional distribution of borrower types observed in a seasoned mortgage pool is endogenously tilted toward lower hazard borrowers. Burnout therefore arises as a purely statistical selection effect rather than a behavioral change in borrower refinancing incentives.
@@ 5.3 Normal Frailty
Strictly speaking, frailty variables in survival models are required to be nonnegative, since they scale the baseline hazard. For this reason the gamma and lognormal distributions are commonly used in applications. The normal distribution does not satisfy this positivity constraint and is therefore not appropriate as a structural frailty model. Nevertheless, it is useful as a simple analytic benchmark illustrating how cross-sectional dispersion generates burnout.
For illustration we therefore also consider the normal distribution.
@@
5.4 Multivariate Frailty
In many mortgage applications borrower heterogeneity is driven by several unobserved factors rather than a single scalar frailty variable. A natural extension is therefore a multivariate frailty model [1], in which the borrower hazard takes the form
$\lambda_{t}(x,y)=f(x)^{\top}\lambda^{0}_{t}(y),$
where $f(x)=(f_{1}(x),\ldots,f_{k}(x))^{\top}$ represents a vector of borrower-specific frailty factors and $\lambda^{0}_{t}(y)$ is a vector of common factor intensities that may depend on the observed loan characteristics $y$. Examples of such latent factors include borrower mobility, refinancing transaction costs, credit constraints, and other behavioral characteristics that are difficult to observe directly but influence the propensity to prepay.
Under this specification the pool hazard becomes
$\bar{\lambda}_{t}(y)=\mathsf{E}_{t}(f(x))^{\top}\lambda^{0}_{t}(y),$
where $\mathsf{E}_{t}(f(x))$ denotes the survival weighted mean frailty vector. The burnout dynamics are therefore governed by the evolution of the survival weighted distribution of the multivariate frailty factors. Applying the deterministic burnout identity yields
$\frac{d}{dt}\,\bar{\lambda}_{t}(y)=\mathsf{E}_{t}(\dot{\lambda}_{t}(x,y))-\lambda^{0}_{t}(y)^{\top}\mathsf{Cov}_{t}(f(x))\,\lambda^{0}_{t}(y),$
where $\mathsf{Cov}_{t}(f(x))$ is the survival weighted covariance matrix of the frailty vector.
Thus in the multivariate setting the burnout effect is driven by the cross-sectional covariance structure of borrower frailties. Directions in factor space with larger dispersion contribute more strongly to the decline of the pool hazard as high frailty borrowers exit the pool. This formulation shows that mortgage burnout can be interpreted as a selection effect operating on a multidimensional latent factor structure for borrower behavior.
@@ 6 Conclusion
The negative variance term provides a structural stochastic analog of burnout: heterogeneity induces a systematic downward drift in the observed hazard through selection effects. The resulting formula may be interpreted as a stochastic continuous time analog of the Price equation, showing that burnout arises from a universal selection mechanism in heterogeneous hazard populations.